write an expression for speed of electromagnetic wave in free space?

Grade:8

1 Answers

9893833800
6 years ago
From electrostatic,1/4 (3.15)e¤=9×10^9 Nm^2C^-2And From magnetism,U¤/4 (3.15)=10^-7 TmA^-1Therefore, u¤e¤ = u¤/4 (3.15). [4 (3.15)e¤] =10^-7 ×1/9×10^9 = 1/(3×10^8)^2We know that, c=3×10^8 ms^-1Therefore,u¤e¤ = 1/c^2OrC^2 =1/e¤e¤Hence, speed of light is inversely equal to the square root of products of permeability and permittivity.i.e. 3×10^8 ms^-1 which is the speed of light i.e. am waves.....